Skip to main content

Milepost 21

Unit Information

Hoopa Valley Tribe
Bureau of Indian Affairs
11860 State Highway 96 P.O. Box 1348 Hoopa California 95546
Hoopa, CA 95546

BIA Shield

Incident Contact

Kate Kramer
Email: kate_kramer@firenet.gov
Phone: 5303242162
Hours: M-F 8am - 8pm, SS 8am - 8 pm

Beginning today, a smaller Type 3 firefighting organization will be in place to complete containment of the fire.  The remaining work is focused on suppression repair.  Portions of the dozer lines remain; they must be treated with water bars and covered with brush and then inspected before considered complete. A wood chipper will be brought in to chip up cut brush; wood chips are spread over areas disturbed by fire line construction to prevent later soil erosion.  An excavator will be used to water-bar and place slash and cut-brush over the dozer lines.  The firefighters will also assist the Hoopa Fire Department as needed with any new starts. 

Weather for the area: Hot and dry weather conditions are expected to continue through the week. Good relative humidity recovery of 50% or higher will occur overnight. Thunderstorms are expected to the north and east of the fire area today; there is a slight chance for storms in the fire area.   Winds of 6-12 mph, with gusts of 25 mph during the warmest part of the afternoon are expected.  

Unless conditions change, this is the final update for the Milepost 21 Fire.  
 

California Interagency Incident Management Team 11 appreciates the support of Hoopa Valley tribal members and the Hoopa Fire Department.  We hope everyone remains healthy during these challenging pandemic times. 

Basic Information

Current as of
Incident TypeWildfire
Date of OriginMonday July 20th, 2020 approx. 12:00 PM
LocationMilepost 21 along Hwy 96, east of highway.
Incident DescriptionWildland Fire
Coordinates41.168 latitude, -123.696 longitude

Current Situation

Total Personnel150
Size100 Acres
Percent of Perimeter Contained90%
Estimated Containment DateSunday July 26th, 2020 approx. 12:00 AM
Fuels Involved

Timber

Significant Events

Interior stump holes continue to smolder, posing no threat to established containment lines.

Outlook

Planned Actions

Operations units will continue to monitor and patrol established containment lines. Suppression repair activates will continue. The management team will brief and prepare the incoming Type III management team. Transfer of command is planned for 0700 hours July 27,2020.

Projected Incident Activity

No fire growth is anticipated.

Current Weather

Weather Concerns

High pressure over region will continue to bring hot and dry conditions over the fire area today with a slight chance of isolated thunderstorms in the late afternoon and early evening hours.

Cumulus buildups may be noticeable to the north and east of the incident and any thunderstorms may contain strong downdrafts that can impact the fire. Expect warmer and drier conditions overnight, but with relative humidity recoveries over the fifty percent.